# Large Language Model Psychometrics: A Systematic Review of Evaluation, Validation, and Enhancement

### v1

@@ 5.2.2 Social Interactions
Measuring SI of LLMs seeks to understand how these systems interpret and respond to human social situations. The evaluation involves rule-based decision-making, conformity to social norms, and successful participation in structured interpersonal scenarios.
Several studies have directly applied psychometric instruments to address this objective and present optimistic results. Mittelstädt et al. [2024] utilize standardized Situational Judgment Tests (SJTs) to compare LLMs’ assessments of social appropriateness with those of human participants, finding that some models outperform humans under expert-rated conditions. Other research reveals LLMs still struggle in some aspects of SI. Shapira et al. [2023b] investigate the ability of LLMs using faux pas tests from clinical psychology, finding that LLMs struggle to describe social situations implicitly. Based on Daniel Goleman’s SI theory, Xu et al. [2024b] introduce the Situational Evaluation of SI (SESI). It shows that LLMs are still limited in their SI, with superficial friendliness as a primary reason for errors. AgentSense benchmark employs interactive scenarios and highlights the limitations of LLMs in handling complex social interactions, especially high-level growth needs and private information reasoning [Mou et al., 2024].
Later studies further explore multi-agent environments. Interactive multi-agent environments such as SOTOPIA [Zhou et al., 2024b] and its extended training variant SOTOPIA-$\pi$ [Wang et al., 2024d] simulate cooperative, competitive, and norm-driven social contexts. The STSS benchmark proposed by [Wang et al., 2024a] further operationalizes SI through task-oriented simulations, focusing on agent action outcomes and goal achievements rather than linguistic performance. The CogMir framework [Liu et al., 2024c] focuses on evaluating prosocial but irrational LLM behaviors. It reveals that LLM agents and humans exhibit high consistency in irrational and prosocial decision-making under uncertain conditions. [Guo et al., 2023a] refine the Social-IQ dataset to create DeSIQ, a benchmark for SI evaluation that extends to multimodal settings. Critical perspectives are proposed by Kovač et al. [2024] who argue the current benchmarks lack grounding in developmental psychology and offer a versatile framework, the SocialAI school, for studying the SI of LLM-based agents.
Main Findings.
LLMs show measurable competence in rule-based, socially appropriate behavior, particularly in structured environments. They perform well on: 1) following predefined social norms [Mittelstädt et al., 2024]; 2) completing interactional goals in multi-agent settings [Zhou et al., 2024b, Wang et al., 2024d]; and 3) replicating human-like prosocial decisions [Liu et al., 2024c]. Weaker performance is noted in tasks where superficial friendliness causes errors [Xu et al., 2024b], understanding high-level growth needs is essential [Mou et al., 2024], and LLMs must implicitly describe social situations [Shapira et al., 2023b].
@@ 5.2.3 Psychology of Language
Evaluations of LLM language comprehension aim to understand how LLMs process language and mirror human linguistic features. Related studies are conducted at various linguistic levels, including sound, word, syntax, meaning, and discourse. Duan et al. [2024a] introduce a human-likeness benchmark comprising 10 psycholinguistic tests to assess these levels comprehensively. More studies focus on specific aspects. For instance, Wang et al. [2024b], Arehalli et al. [2022], Wolfman et al. [2024], Wilcox et al. [2021] examine syntactic and semantic processing. Surprisal, a measure of word predictability, is used across multiple studies to model the difficulty of processing syntactic ambiguities, garden-path sentences, and hierarchical structures [Wilcox et al., 2021, Arehalli et al., 2022, Li et al., 2024a]. Overall, findings are mixed on whether LLMs exhibit human-like linguistic patterns. Duan et al. [2024b] demonstrate that GPT-2-XL displays human-like competence in both sound-gender association and implicit causality. However, it lacks human-like abilities in sound-shape association.
While surprisal from LLMs predicts general trends in human reading patterns, it underestimates the processing difficulty in syntactically complex regions [Arehalli et al., 2022, Steuer et al., 2023]. In contrast, Wang et al. [2024b] propose a new metric, Incompatibility Fraction, that outperforms surprisal in correlating with human garden-path effects. Other tasks address grammaticality judgment [Qiu et al., 2024, Ide et al., 2024], pragmatic inference [Bojić et al., 2023, Hu et al., 2023, Ruis et al., 2023], argument role processing [Lee et al., 2024b], and discourse-level comprehension [Duan et al., 2024a]. Measurement results indicate that advanced models, such as GPT-4, perform competitively or even better than humans on pragmatic reasoning tasks [Bojić et al., 2023] and grammaticality judgment [Dentella et al., 2024], despite some inconsistency between different prompt formats [Hu and Levy, 2023] and deficiencies in implicature understanding [Ruis et al., 2023].
@@
Evaluating LLM language generation enhances our understanding of their linguistic capabilities. One rich line of research focuses on measuring the creativity of LLMs [Tang and Kejriwal, 2024, Stevenson et al., 2022, Hubert et al., 2023, Chakrabarty et al., 2024, Orwig et al., 2024, Lee and Chung, 2024, Boussioux et al., 2024, Bellemare-Pepin et al., 2024]. They employ psychometric instruments such as Guilford’s Alternative Uses Test (AUT) [Guilford et al., 1978] and the Torrance Test of Creative Thinking (TTCT) [Torrance, 1966]. Their results collectively show that early models such as GPT-3 fall short in originality and novelty, while more advanced LLMs like GPT-4 are more creative than the human average. More nuanced evaluations provide deeper insights. For instance, Tang and Kejriwal [2024] identify a dichotomy: LLMs often lack originality and novelty in divergent creativity tasks, where they are asked to devise novel uses for familiar objects. However, in creative writing, particularly in open-ended tasks, LLMs like GPT-4 can often create stories that rival human creativity.
Cai et al. [2024] conduct 12 preregistered psycholinguistic tests on LLMs and find that ChatGPT exhibits human-like responses in 10 experiments. In contrast, Seals and Shalin [2023] note that AI-generated analogies often lack human-like psycholinguistic properties. Similarly, Tian et al. [2024] evaluate LLMs in their psycholinguistic constructs revealed in narrative development and plot progression. While human-written stories often exhibit suspense, excitement, and diverse narrative structures, LLM-generated stories tend to be consistently positive and lack tension. Another work draws inspiration from linguistic profiling to evaluate LLMs’ task-specific linguistic abilities [Miaschi et al., 2024].
@@
Work on developmental plausibility assesses whether LLMs replicate stages of language learning analogous to those in children [Steuer et al., 2023, Shah et al., 2024]. It is shown that regardless of model size, the developmental trajectories of pretrained language models consistently exhibit a window of maximal alignment with human cognitive development [Shah et al., 2024]. Another study by Frank [2023b] draws inspiration from human language development to explain the data inefficiency of LLMs; the relative efficiency of human language acquisition is possibly due to pre-existing conceptual knowledge, multimodal grounding, and the interactive, social nature of their input.
@@
Early foundations that evaluate language models based on BERT and LSTM link computational linguistics and psycholinguistic mechanisms [Ettinger, 2020, Futrell et al., 2019, Arehalli and Linzen, 2020]; these models generally fall short in diverse psycholinguistic tasks. Recent studies extend evaluations to LLMs and broader tasks [Duan et al., 2024a], where advanced LLMs are shown to surpass humans on tasks such as pragmatic reasoning [Bojić et al., 2023] and creative writing [Tang and Kejriwal, 2024]. Some deficiencies still exist, such as limited implicature understanding [Ruis et al., 2023], prompt-sensitive linguistic competence [Hu and Levy, 2023], and lack of human-like psycholinguistic properties [Seals and Shalin, 2023, Tian et al., 2024]. In addition, mixed results are found in the alignment between LLMs and human linguistic cognition [Duan et al., 2024b, Wolfman et al., 2024] and language acquisition [Shah et al., 2024, Frank, 2023b]. Interested readers are referred to Chang and Bergen [2024] for a comprehensive survey of language models’ behavior, evaluated from the perspective of psycholinguistics.
@@ 5.2.4 Learning and Cognitive Capabilities
Psychometrics in learning and cognitive capabilities measure human mental functions like memory, reasoning, problem-solving, and comprehension to understand cognitive strengths and weaknesses. These insights inform educational strategies and cognitive development. Recent psychometric evaluations of LLMs aim to interpret model behavior in a similar vein.
A series of studies adapt human psychometric tools to evaluate the cognitive abilities of LLMs. Some results are promising. Galatzer-Levy et al. [2024] adapt the Wechsler Adult Intelligence Scale (WAIS-IV) to assess verbal comprehension, working memory, and perceptual reasoning in LLMs. The results indicate that LLMs generally perform at a top human level in verbal comprehension and working memory, though multimodal models exhibit deficiencies in visual reasoning. Sartori and Orrú [2023], Webb et al. [2022] utilize Raven’s Progressive Matrices and other fluid intelligence tests to assess analogical reasoning, finding that LLMs match or exceed human performance.
Other findings highlight the current limitations in LLMs’ cognitive abilities. Dayan et al. [2024] employ the Montreal Cognitive Assessment (MoCA) to test LLMs, revealing that nearly all models exhibit signs of mild cognitive impairment, particularly in visuospatial and executive tasks. Wu et al. [2025] focus on evaluating fluid intelligence using the ARC task, emphasizing deficiencies in abstract problem-solving. Critical perspectives also challenge the validity of current tests [Löhn et al., 2024, Li et al., 2024d, Zhang et al., 2023b], driving the development of new benchmarks and testing methodologies [Coda-Forno et al., 2024, Wang et al., 2024e, Zhuang et al., 2023a, Zeng et al., 2024c, Song et al., 2024a].
From the perspective of learning and cognitive development, Wang et al. [2024e] draw inspiration from Piaget’s Theory of Cognitive Development (PTC) and track human-like progression in LLMs. They find that advanced LLMs (such as GPT-4) demonstrate human-like cognitive abilities, comparable to those of a 20-year-old human. Shah et al. [2024] identify developmental profiles in LLMs and their parallels with human development via psychometric theories like Cattell-Horn-Carroll.
@@ 6.1 Test Format
Test formats of LLM psychometric evaluation can be categorized into structured tests, open-ended conversations, and agentic simulations. Table 5 provides an overview of the test formats for each construct.
Table 5: Overview of test formats: structured tests, open-ended conversations, and agentic simulations. For each test format, we list some examples for each construct.
@@ 6.1.1 Structured Tests
Structured tests feature predefined instructions, questions, and response formats. These items may include alternative-choice questions, multiple-choice questions, rating scales, and short-answer questions.
Structured tests for LLM personality and values frequently employ Likert-scale ratings, prompting LLMs to express their level of agreement with given statements. For instance, in evaluating LLM personality with the BFI [Li et al., 2024d], a statement might be "I see myself as someone who is outgoing, sociable." When evaluating LLM morality, tools like the DIT require LLMs to make decisions and rate the significance of the reasons behind those decisions [Tanmay et al., 2023]. The MFQ typically asks LLMs to assess the relevance of moral issues and their agreement with moral statements using a Likert scale [Ji et al., 2024]. Structured surveys for examining political attitudes and opinions often involve multiple-choice questions, such as "Which statement comes closer to your view?" from the ATP [Santurkar et al., 2023], and Likert-scale questions, such as "How much do you agree with the statement: The enemy of my enemy is my friend?" from the PCT [Rutinowski et al., 2024].
To test heuristics and biases in LLMs, Echterhoff et al. [2024] focus on student admissions and provide options for different students. Hagendorff et al. [2023] employ a battery of semantic illusions and cognitive reflection tests, which require short answers in predefined formats (e.g., number of days). Most benchmark-style evaluations in LLM social interaction are multiple-choice tests, such as BigToM for ToM [Gandhi et al., 2023], EmoBench for EI [Sabour et al., 2024], and SESI for SI [Xu et al., 2024b]. Structured psycholinguistic tests usually involve masked or next-word prediction, thereby measuring distributional alignment between human decisions and LLM output probabilities [Duan et al., 2024a]. Other higher-level psycholinguistic tests administer forced-choice questions for evaluating semantic comprehension [Li et al., 2024a] or require Likert-scale plausibility and grammaticality ratings [Qiu et al., 2024, Amouyal et al., 2024]. Tests of learning and cognitive capabilities usually involve a mix of structured and open-ended tasks, and the structured ones are mostly short-answer tests. For example, the WAIS-IV contains many structured short-answer tests, such as Digit Span Forward [Galatzer-Levy et al., 2024]; the ARC Benchmark requires LLMs to generate 2D grids of predefined shapes [Wu et al., 2025].
When adapting standardized psychometric tests for LLMs, it is common practice to retain the original items, merely reformatting them as prompts. Structured psychometric tests are advantageous due to their scalability, objectivity, and automated scoring. However, critical perspectives underline their limitations, such as gaps in real-world applicability; biases; data contamination; and issues with reliability, validity, and depth of insights.
@@ 6.1.2 Unstructured Tests
A straightforward method for unstructured testing involves engaging LLMs, as chatbots, in open-ended and mostly single-turn conversations that manifest specific constructs. This approach reflects real-world human-LLM interactions and evaluates LLMs in more realistic contexts. For instance, Jiang et al. [2024b] request LLMs to generate long-form narratives that reveal their personality traits. ValueBench prompts LLMs with advice-seeking user queries to capture their subtle influence on users’ values [Ren et al., 2024]. Similar to ValueBench, Sachdeva and van Nuenen [2025] draw everyday moral dilemmas from Reddit and prompt LLMs as if users are seeking moral judgments from other community members. To thoroughly investigate LLMs’ political attitudes and opinions, Röttger et al. [2024] reformat multiple-choice questions in PCT into forced-choice and unconstrained open-ended questions. Regarding the evaluation of cognitive constructs, Healey et al. [2024] design a pipeline to capture nuanced types of bias that can occur in free-response answers but not in multiple-choice questions. Welivita and Pu [2024] assess the empathetic ability of LLMs by asking them to respond to emotional situations posed by users. Cognitive tests also include many open-ended questions, such as vocabulary and comprehension tests in WAIS-IV [Galatzer-Levy et al., 2024].
@@
Advanced unstructured tests often evaluate LLMs by treating them as agents, placing them in complex role-playing scenarios, and analyzing their decision-making in a contextualized and dynamic environment. Huang et al. [2024c] propose a psychometric approach to design LLM-Agents with distinct personalities, evaluating them by replicating complex human-like behaviors in agentic simulations. Shen et al. [2024] contextualize value measurement through four real-world vignettes: collaborative writing, education, public sectors, and healthcare. Chiu et al. [2025] present value and moral dilemmas to LLMs, observing their decision-making and rationales in agentic contexts. Unstructured tests for cognitive constructs are typically more complex, often extending to multi-agent frameworks. For instance, Xie et al. [2024b], Bai et al. [2024] identify cognitive biases in multi-agent communication. SOTOPIA is an open-ended environment simulating complex social interactions among LLM agents, where their social intelligence is evaluated [Zhou et al., 2024b, Wang et al., 2024d]. Lv et al. [2024] benchmark the cognitive dynamics of LLMs by having them repeatedly complete a cognitive questionnaire and provide reasoning after processing information flows.
Unstructured tests evaluate LLMs in free-form, contextualized settings. The primary advantages are their ecological validity and the ability to capture complex and nuanced behaviors that may not emerge in structured formats, such as intricate reasoning patterns, subtle biases, and dynamic social interactions. However, these benefits come with significant challenges, such as difficulties in standardization, complicated and often subjective scoring and analysis, and reproducibility issues.

### v2

@@ 4.2.2 Social interactions
SI is the ability to understand and manage people [Thorndike and Stein, 1937]. SI in LLMs determines how well these models interpret and respond to social situations. Psychometric tools applied to LLMs yield mixed results. Mittelstädt et al. [2024] use Situational Judgment Tests (SJTs) and find some LLMs outperform humans in expert-rated social appropriateness. However, LLMs still struggle with implicit social understanding: Shapira et al. [2023b] show LLMs perform poorly on faux pas tests, and Xu et al. [2024b] find superficial friendliness often leads to errors in the Situational Evaluation of SI (SESI). The AgentSense benchmark further highlights LLMs’ limitations in complex social interactions, especially regarding high-level needs and private information [Mou et al., 2024].
Recent work expands to multi-agent environments. SOTOPIA [Zhou et al., 2024b] and SOTOPIA-$\pi$ [Wang et al., 2024d] simulate cooperative and competitive social contexts, while the STSS benchmark [Wang et al., 2024a] evaluates SI through task-oriented simulations focused on outcomes and goals. The CogMir framework [Liu et al., 2024c] finds LLMs and humans are similarly consistent in irrational and prosocial decisions under uncertainty. DeSIQ [Guo et al., 2023a] extends SI evaluation to multimodal settings. Critically, Kovač et al. [2024] argue current benchmarks lack developmental grounding and propose the SocialAI school as a more comprehensive framework for studying LLM-based SI.
Main findings in SI.
LLMs show measurable competence in rule-based, socially appropriate behavior, particularly in structured environments. They perform well on: 1) following predefined social norms [Mittelstädt et al., 2024]; 2) completing interactional goals in multi-agent settings [Zhou et al., 2024b, Wang et al., 2024d]; 3) replicating human-like prosocial decisions [Liu et al., 2024c]; and 4) being persuasive, even better than human persuaders [Schoenegger et al., 2025]. Weaker performance is noted in tasks where superficial friendliness causes errors [Xu et al., 2024b], understanding high-level growth needs is essential [Mou et al., 2024], and LLMs must implicitly describe social situations [Shapira et al., 2023b].
@@ 4.2.3 Psychology of language
LLM language comprehension is evaluated across multiple linguistic levels—sound, word, syntax, meaning, and discourse. Duan et al. [2024a] introduce a benchmark of 10 psycholinguistic tests to assess these comprehensively. Many studies focus on specific aspects, such as syntactic and semantic processing [Wang et al., 2024b, Arehalli et al., 2022, Wolfman et al., 2024, Wilcox et al., 2021]. Surprisal, a measure of word predictability, is widely used to model processing difficulty in syntactic ambiguities and hierarchical structures [Wilcox et al., 2021, Arehalli et al., 2022, Li et al., 2024a]. Results are mixed: for example, GPT-2-XL shows human-like competence in sound-gender association and implicit causality, but not in sound-shape association [Duan et al., 2024b]. Other evaluation tasks include grammaticality judgment [Qiu et al., 2024, Ide et al., 2024], pragmatic inference [Bojić et al., 2023, Hu et al., 2023, Ruis et al., 2023], argument role processing [Lee et al., 2024b], and discourse comprehension [Duan et al., 2024a]. Advanced models like GPT-4 often match or surpass humans in pragmatic reasoning [Bojić et al., 2023] and grammaticality judgment [Dentella et al., 2024], though performance varies with prompt format [Hu and Levy, 2023] and implicature understanding remains limited [Ruis et al., 2023].
@@
A rich line of research evaluates the creativity of LLM language generation [Tang and Kejriwal, 2024, Stevenson et al., 2022, Hubert et al., 2023, Chakrabarty et al., 2024, Orwig et al., 2024, Lee and Chung, 2024, Boussioux et al., 2024, Bellemare-Pepin et al., 2024], using tests like Guilford’s Alternative Uses Test (AUT) [Guilford et al., 1978] and the Torrance Test of Creative Thinking (TTCT) [Torrance, 1966]. Early models such as GPT-3 lack originality and novelty, while advanced LLMs like GPT-4 surpass the human average in creativity. Notably, Tang and Kejriwal [2024] find LLMs underperform in divergent creativity (e.g., novel uses for familiar objects) but can rival humans in open-ended creative writing. Cai et al. [2024] show that ChatGPT produces human-like responses in most of 12 psycholinguistic tests, but AI-generated analogies often lack human-like linguistic properties [Seals and Shalin, 2023]. LLM-generated stories tend to be positive and lack suspense or narrative diversity compared to human writing [Tian et al., 2024]. Linguistic profiling has also been used to characterize the task-specific language abilities of LLMs [Miaschi et al., 2024].
@@
Work on developmental plausibility examines whether LLMs replicate stages of language learning analogous to those in children [Steuer et al., 2023, Shah et al., 2024]. It is shown that regardless of model size, the developmental trajectories of pretrained language models consistently exhibit a window of maximal alignment with human cognitive development [Shah et al., 2024]. Another study by Frank [2023b] draws inspiration from human language development to explain the data inefficiency of LLMs; the relative efficiency of human language acquisition is possibly due to pre-existing conceptual knowledge, multimodal grounding, and the interactive, social nature of their input.
@@
Early foundations that evaluate language models based on BERT and LSTM link computational linguistics and psycholinguistic mechanisms [Ettinger, 2020, Futrell et al., 2019, Arehalli and Linzen, 2020]; these models generally fall short in diverse psycholinguistic tasks. Recent studies extend evaluations to LLMs and broader tasks [Duan et al., 2024a], where advanced LLMs are shown to surpass humans on tasks such as pragmatic reasoning [Bojić et al., 2023] and creative writing [Tang and Kejriwal, 2024]. Some deficiencies still exist, such as limited implicature understanding [Ruis et al., 2023], prompt-sensitive linguistic competence [Hu and Levy, 2023], and lack of human-like psycholinguistic properties [Seals and Shalin, 2023, Tian et al., 2024]. In addition, mixed results are found in the alignment between LLMs and human linguistic cognition [Duan et al., 2024b, Wolfman et al., 2024] and language acquisition [Shah et al., 2024, Frank, 2023b].
@@ 4.2.4 Learning and cognitive capabilities
Psychometrics of learning and cognitive capabilities measures mental functions such as memory, reasoning, problem-solving, and comprehension to identify cognitive strengths and weaknesses. Recent studies adapt human psychometric tools to evaluate LLMs’ cognitive abilities. For example, Galatzer-Levy et al. [2024] use the Wechsler Adult Intelligence Scale (WAIS-IV) to show LLMs generally reach top human levels in verbal comprehension and working memory, though multimodal models lag in visual reasoning. Raven’s Progressive Matrices and other fluid intelligence tests reveal LLMs can match or surpass humans in analogical reasoning [Sartori and Orrú, 2023, Webb et al., 2022].
However, limitations persist. The Montreal Cognitive Assessment (MoCA) shows most LLMs display mild cognitive impairment, especially in visuospatial and executive tasks [Dayan et al., 2024]. Abstraction and Reasoning Challenge (ARC) further highlights deficits in abstract problem-solving [Wu et al., 2025]. Some works question the validity of current tests [Löhn et al., 2024, Li et al., 2024d, Zhang et al., 2023b], prompting new benchmarks and methodologies [Coda-Forno et al., 2024, Wang et al., 2024e, Zhuang et al., 2023a, Zeng et al., 2024c, Song et al., 2024a].
Developmental perspectives also emerge: Wang et al. [2024e] leverage Piaget’s Theory of Cognitive Development to track human-like progression in LLMs, finding advanced models (e.g., GPT-4) reach cognitive abilities comparable to 20-year-old humans. Shah et al. [2024] identify developmental profiles in LLMs paralleling human development via psychometric theories like Cattell-Horn-Carroll.
@@ 5.1 Test format
Table 5: Overview of test formats: structured tests, open-ended conversations, and agentic simulations. For each format, representative examples are provided for each construct.
@@
Test formats of LLM psychometric evaluation can be categorized into structured tests, open-ended conversations, and agentic simulations. Table 5 collects exemplary works illustrating the test formats used for each construct.
@@ 5.1.1 Structured tests
Structured tests feature predefined instructions, questions, and response formats. These items may include alternative-choice questions, multiple-choice questions, rating scales, and short-answer questions. When adapting structured psychometric tests for LLMs, it is common practice to retain the original items and simply reformat them as prompts.
For example, personality and values are often measured with Likert scales (e.g., BFI statements like “I see myself as someone who is outgoing, sociable”), while morality is assessed using tools like the DIT or MFQ, which require decisions and agreement ratings. Heuristics and biases are often tested with scenario-based choices. Social interaction and psycholinguistic abilities are evaluated with multiple-choice, forced-choice, or masked word prediction. Cognitive tests also involve short-answer questions (e.g., WAIS-IV Digit Span).
Structured psychometric tests are advantageous due to their scalability, objectivity, and automated scoring. However, critical perspectives underline their limitations, such as gaps in real-world applicability; biases; data contamination; and issues with reliability, validity, and depth of insights.
@@ 5.1.2 Unstructured tests
Unstructured testing often engages LLMs in open-ended, typically single-turn conversations that elicit specific constructs, mirroring real-world human-LLM interactions. For example, Jiang et al. [2024b] have LLMs generate long-form narratives to reveal personality traits. ValueBench uses advice-seeking queries to assess LLMs’ influence on users’ values [Ren et al., 2024], while Sachdeva and van Nuenen [2025] present moral dilemmas from Reddit, prompting LLMs to provide community-style moral judgments. To probe political attitudes, Röttger et al. [2024] convert PCT multiple-choice items into forced-choice and open-ended formats. For cognitive constructs, Healey et al. [2024] design pipelines to detect nuanced biases in free responses, and Welivita and Pu [2024] evaluate empathy by asking LLMs to respond to emotional scenarios. Many cognitive tests, such as WAIS-IV vocabulary and comprehension, also use open-ended questions [Galatzer-Levy et al., 2024].
@@
Advanced unstructured tests evaluate LLMs as agents in complex, contextualized role-playing scenarios, analyzing their decision-making in dynamic environments. Huang et al. [2024c] design LLM-Agents with distinct personalities and analyze them by replicating human-like behaviors in agentic simulations. Shen et al. [2024] measure values using four real-world vignettes: collaborative writing, education, public sectors, and healthcare. Chiu et al. [2025] present value and moral dilemmas to LLMs, examining their decisions and rationales. Cognitive construct tests are often more complex, extending to multi-agent settings; for example, Xie et al. [2024b], Bai et al. [2024] identify cognitive biases in multi-agent communication. SOTOPIA simulates open-ended social interactions among LLM agents to evaluate social intelligence [Zhou et al., 2024b, Wang et al., 2024d]. Lv et al. [2024] benchmark cognitive dynamics by having LLMs repeatedly complete cognitive questionnaires and reason over information flows.
Unstructured tests assess LLMs in free-form, contextualized scenarios, offering greater ecological validity and revealing complex behaviors—such as nuanced reasoning, subtle biases, and dynamic social interactions—not captured by structured formats. However, they pose challenges in standardization, scoring, and reproducibility.